Problem Requirements 1. Calculate Interest income for periods indicated below, using the terms listed.

Note Receivable Terms

Issue Date January 1, 2019

Maturity Date (6 Months) June 30, 2019

Principal $ 12,000

Interest Rate - Annual Percentage Rate (APR) 8.0%

Calculate Interest Income for Questions 1 - 3 (Show calculations)

1. First three months of 2019

2. January 2019 (One Month)

3. Total Interest Income to Maturity in 2019

4. Indicate below the total amount of cash the company will collect from the Note Receivable in 2019, including interest and principal.
Answer #1

1) Interest for first three months = => $12000*8%*3/12=$240.

2)Interest for January Month = $12000*8%*1/12 = $80.

3)Total amount of cash = Principal + Interest.

=> $12,000 + ($12,000 * 8% * 6/12) = $12000+$480

=> $12,480.

IF YOU HAVE ANY DOUBTS PLEASE COMMENT ON THE ISSUE.